40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 10 23:14:01 2011, kiwamu, Update, SUS, ETMX_OPLEV : He-Ne laser dead ? 
    Reply  Wed May 11 11:05:45 2011, steve, Update, SUS, ETMX_OPLEV : He-Ne replaced P1070672.JPG
       Reply  Wed May 11 21:58:30 2011, rana, Update, SUS, ETMX_OPLEV : He-Ne replaced 
Message ID: 4693     Entry time: Wed May 11 21:58:30 2011     In reply to: 4686
Author: rana 
Type: Update 
Category: SUS 
Subject: ETMX_OPLEV : He-Ne replaced 

Quote:

Quote:

I found that a He-Ne laser which has been used for ETMX_OPLEV was NOT giving the light.

Since I didn't find the switch key for it I have no idea if the laser is simply off or dead.

 The dead laser was replaced by new JDSU 1104P of  2.6mW. The return beam is big ~5 mm diameter of 0.3 mW,  1400 counts

 Whenever replacing any Oplev laser, please also put into the ELOG when it was installed so that we have an electronic record of the laser lifetime.

ELOG V3.1.3-